OpenStudy (anonymous):

Solve a system of 4 equations: w+x+y+z=1 2w+x-y-z=4 w-x-2y+2z=2 3w+2x+y+z=7 The answer is supposed to be 7. How do I get there?

OpenStudy (anonymous):

Oops. not 7

OpenStudy (zarkon):

just 7?...interesting

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

w is supposed to equal 7

OpenStudy (anonymous):

canceling out variables.

OpenStudy (anonymous):

I need to find w z y & z

OpenStudy (anonymous):

i mean: w x y and z

OpenStudy (anonymous):

7 -8 4.25 -2.25

OpenStudy (anonymous):

Whoa. how did you get to the answers that quickly?

OpenStudy (anonymous):

use excel and Minverse, (array result requires you hold command when press enter), then MMult with the solution vector. Your TI-83 can also do matrices. Don't bother with elimination, it takes too long, there's better things to do.

OpenStudy (anonymous):

I need the steps. I'm working on a project.

OpenStudy (anonymous):

do you know how to build a matrix from a set of equations?

OpenStudy (anonymous):

@Tonks No I don't.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Ok, take the co-efficient of each term, if there isn't one, then it is 1: 2w+x-y-z=4 becomes 2, 1, -1, -1 in a coefficient matrix and the 4 goes in your solution vector (a matrix that is only one column wide is called a vector) w+x+y+z=1 2w+x-y-z=4 w-x-2y+2z=2 3w+2x+y+z=7 becomes: 1, 1, 1, 1 2, 1, -1, -1 1, -1, -2, 2 3, 2, 1, 1 for your coefficient matrix your solution vector is: 1 4 2 7

OpenStudy (anonymous):

So far so good?

OpenStudy (anonymous):

Exact solutions from Mathematica:\[\left\{w = 7,x = -8,y = \frac{17}{4},z = -\frac{9}{4}\right\} \]

OpenStudy (anonymous):

That might be a little complicated for the level I am on. I was thinking more in the direction of adding them together to eliminate variables.

OpenStudy (anonymous):

oof, the problem with using elimination is that there are so many opportunities for errors, while there are only opportunities for entry errors with the matrix method.

OpenStudy (anonymous):

That's what my teacher wants us to do...

OpenStudy (anonymous):

The following is a matrix solution using Mathematica. LinearSolve finds an x which solves the matrix equation m.x==b. \[\text{LinearSolve}\left[\left( \begin{array}{cccc} 1 & 1 & 1 & 1 \\ 2 & 1 & -1 & -1 \\ 1 & -1 & -2 & 2 \\ 3 & 2 & 1 & 1 \\ \end{array} \right),\left( \begin{array}{c} 1 \\ 4 \\ 2 \\ 7 \\ \end{array} \right)\right]\to \left\{7,-8,\frac{17}{4},-\frac{9}{4}\right\} \]

OpenStudy (anonymous):

Well I have to go. Thanks for your help!